# Qubit encoding for a mixture of localized functions

### v2

@@ I Introduction
To this end, we design the main encoding technique by starting from the probabilistic operation [10, 11] for an LC of discrete Lorentzian functions (LFs), to which we apply the quantum amplitude amplification (QAA) technique [12, 13] to render the encoding deterministic. As is demonstrated later, expanding a target function in LFs is favorable for achieving efficient circuit implementation. Our encoding techniques possess applicability to diverse fields of quantum computation. Amongst them, quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] is a promising one, where an initial state can be constructed from one-electron molecular orbitals (MOs) [25, 26], typically expressed as LCs of localized orbitals. We provide resource estimation for such calculations. To demonstrate the efficiency and practical usefulness of our scheme, we perform it on the real quantum computers commercially provided by IBM. Our encoding scheme is for a dense quantum state; that is, it treats a target function that has basically a nonzero amplitude of each computational basis. For encoding schemes for a sparse quantum state, see, e.g., Refs [27, 28, 29] and references therein.
@@ II.1.1 Probabilistic operation of an LC of unitaries and amplitude reduction
In this study, we define another circuit $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ for the probabilistic operation of $\mathcal{A}.$ This circuit differs from $\mathcal{C}_{\mathrm{prob}}$ only in that the new one uses an extra ancilla, as depicted in Fig. 1(b). The extra ancilla is accompanied by the $y$ rotation gate $R(2\theta_{\mathrm{AR}})=\exp(-i\theta_{\mathrm{AR}}\sigma_{y})$ with an angle parameter $\theta_{\mathrm{AR}}.$ Although $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ involves $\mathcal{C}_{\mathrm{prob}}$ as its part separated from the extra ancilla, they will be coupled for QAA later. The state $|\Psi_{\mathrm{AR}}\rangle$ of the whole system immediately before the measurement is obviously the tensor product of $|\Psi\rangle$ in Eq. (1) and the extra-ancilla state $\cos\theta_{\mathrm{AR}}|0\rangle+\sin\theta_{\mathrm{AR}}|1\rangle.$ We regard only $|0\rangle_{n_{\mathrm{A}}}\otimes|0\rangle$ to be the success state, even though $|0\rangle_{n_{\mathrm{A}}}\otimes|1\rangle$ also leads to $|\psi_{\mathrm{lc}}\rangle.$ The success probability is thus given by
@@
Figure 1: Each of circuits (a) $\mathcal{C}_{\mathrm{prob}}$ and (b) $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ is for probabilistic operation of an LC $\mathcal{A}$ of unitaries. The explicit construction of the central part $\mathcal{C}$ in $\mathcal{C}_{\mathrm{prob}}$ is described in the original paper [11]. $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ implements the amplitude reduction for the success state by introducing an extra ancilla in addition to those in $\mathcal{C}_{\mathrm{prob}}.$ (c) Each circle represents the weight $W_{\mathrm{f}}$ of failure state in the resultant state of QARA with a QAE error $\Delta w$ as a function of the true weight $w$ of success state in the original state. Each cross represents $W_{\mathrm{f}}$ for QAA with a QAE error. Horizontal lines represent the approximate maximum errors in Eq. (6) for the values of $\Delta w/w.$
@@ II.1.2 Determinization via amplitude reduction and amplification
These observations indicate that $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ with $\theta_{\mathrm{AR,opt}}$ and the subsequent QAA render the nonunitary operation of $\mathcal{A}$ deterministic regardless of $w.$ We refer to this treatment as the determinization of the probabilistic operation and denote it by the quantum amplitude reduction and amplification (QARA) technique. The determinized circuit no longer needs measurements. Since $w$ is on the order of $1/n_{\mathrm{loc}},$ we have $m_{\mathrm{opt}}=\mathcal{O}(\sqrt{n_{\mathrm{loc}}})$ from Eq. (4). In particular, for a case where $1/4\leq w<1,$ we have $m_{\mathrm{opt}}=1.$ Nishi et al. [40] have proposed a similar technique for determinizing the QAA for probabilistic imaginary-time evolution (PITE)[18, 20, 41, 22] steps. The quadratic speedup of PITE [42] with determinization will be a good scheme for nonvariational energy minimization on fault-tolerant quantum computers.
@@ II.1.3 Error analysis for inaccurate QAE
Let us consider a practical situation where the observer obtains an erroneous estimation $w^{(\mathrm{est})}=w+\Delta w$ of the weight $w$ of the original success state. The circuit parameter $\theta_{\mathrm{AR,opt}}$ for amplitude reduction and the repetition number $m_{\mathrm{opt}}$ for amplitude amplification can now involve errors. We demonstrate in Appendix A that QARA allows us to derive a simple relation between the error in the QAE process and that in the determinization. When we approach the limit of $w\to 0$ while keeping $\Delta w/w$ finite, an approximate maximum of the weight of the failure state after the completion of the QARA process is given by
@@
Figure 1(c) shows the weight $W_{\mathrm{f}}$ of the failure state after the QARA process for $\Delta w/w=0.1,0.04,$ and $0.01$ as a function of $w.$ $W_{\mathrm{f}}$ after the QAA process (without amplitude reduction) is also shown in the figure. While the values of $W_{\mathrm{f}}$ from QAA are seen to scatter for each of $\Delta w/w,$ those from QARA exhibit much smaller variations over the wide range of $w.$ Furthermore, $\epsilon_{\mathrm{QARA}}$ approximates $W_{\mathrm{f}}$ well. These observations suggest that QARA is an error-controllable method for determinizing a probabilistic operation.
@@ II.2.1 Setup
Let us consider a case where an $n_{q}$-qubit data register is available and we are provided with $n_{\mathrm{loc}}$ real functions $\{f_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}$ localized at the origin as an expansion basis set in one-dimensional space. For the equidistant grid points $x_{j}\equiv j\Delta x\ (j=0,\dots,2^{n_{q}}-1)$ of a spacing $\Delta x$ on a range $[0,N\Delta x],$ we want to encode a normalized LC of the displaced basis functions on the data register as
@@
where $N\equiv 2^{n_{q}}.$ $|j\rangle_{n_{q}}$ is the computational basis for the data register. $k_{\mathrm{c}\ell}$ is the integer coordinate $(0\leq k_{\mathrm{c}\ell}\leq N-1)$ of the center of the displaced $\ell$th basis function. $d_{\ell}$ is the real coefficient for the LC. We assume that the basis functions are normalized over the range, that is, $\sum_{j=0}^{N-1}f_{\ell}(x_{j})^{2}=1$ for each $\ell,$ and the displaced ones are not necessarily orthogonal to each other. We further assume that the $n_{q}$-qubit unitary for generating each basis function centered at the origin is known: $U_{\mathrm{orig}}^{(\ell)}|0\rangle_{n_{q}}=\sum_{j=0}^{N-1}f_{\ell}(j\Delta x)|j\rangle_{n_{q}}$ for each $\ell.$
@@ II.2.2 Displaced single basis function
for each $\ell.$ To this end, we define the phase shift gate $U_{\mathrm{shift}}(k)$ for an integer $k$ that acts on a computational basis diagonally as
@@
This gate can be implemented as separate single-qubit gates, as explained in Appendix B. With this and the quantum Fourier transform (QFT) [44], the operator
$T(k)\equiv\mathrm{QFT}\cdot U_{\mathrm{shift}}(k)\cdot\mathrm{QFT}^{\dagger}$ (10)
@@ II.2.3 LC of displaced basis functions
$\widetilde{U}^{(\ell)}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathrm{QFT}^{\dagger}\cdot U_{\mathrm{orig}}^{(\ell)}$ (11)
with $U_{\mathrm{shift}}^{(\ell)}\equiv U_{\mathrm{shift}}(k_{\mathrm{c}\ell})$ for each $\ell.$ With the LC $\mathcal{L}\equiv\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}\widetilde{U}^{(\ell)}$ of them, the desired state in Eq. (7) is written as $|\psi_{\mathrm{lc}}\rangle=\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}|\widetilde{f}_{\ell}\rangle=\mathrm{QFT}\cdot\mathcal{L}|0\rangle_{n_{q}}.$ $\mathcal{L},$ which is nonunitary in general, admits the probabilistic implementation as a special case of $\mathcal{C}_{\mathrm{prob}}$ in Fig. 1(a). Specifically, the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the data register and
@@
ancillae encodes $|\psi_{\mathrm{lc}}\rangle$ probabilistically. The central part, $\mathcal{C}_{\mathcal{L}},$ implements $\mathcal{L}$ probabilistically by recognizing the $n_{\mathrm{A}}$-bit integer $\ell$ specified by an ancillary state $|\ell\rangle_{n_{\mathrm{A}}}.$ For details, see the original paper [11], which describes the implementation of a generic LC of unitaries. It is noted here that a slight modification to the circuit in the original paper exists in $\mathcal{C}_{\mathcal{L}}.$ That is, since $\mathrm{QFT}^{\dagger}$ in $\widetilde{U}^{(\ell)}$ does not refer to $\ell$ [see Eq. (11)], the controlled $\mathrm{QFT}^{\dagger}$ gates have been replaced by the uncontrolled $\mathrm{QFT}^{\dagger}$ gate, as seen in Fig. 2(b). Since the $\mathrm{QFT}$ operation after $\mathcal{C}_{\mathcal{L}}$ does not act on the ancillae, it is also possible to perform the measurement, observe the success state, and then operate $\mathrm{QFT}$ on the data register.
Let us estimate the depth of $\mathcal{C}_{\mathrm{lc}}$ by assuming that of $U_{\mathrm{orig}}^{(\ell)}$ is independent of $\ell.$ Since $\mathrm{depth}(\mathrm{QFT})=\mathcal{O}(n_{q})$ [45, 46] and $\mathrm{depth}(U_{\mathrm{shift}}^{(\ell)})=\mathcal{O}(1)$ for each $\ell,$ we find
@@ II.3.2 Discrete Lorentzian functions
$|L;a,0\rangle_{n_{q}}$ is symmetric around $j=N/2.$ As demonstrated in Appendix G.1, the SF and LF states with a common decay rate at the origin are related with each other via QFT: $|L;a,0\rangle_{n_{q}}=\mathrm{QFT}|S;a,0\rangle_{n_{q}}.$ This means that the unitary
$U^{(\mathrm{L})}\equiv\mathrm{QFT}\cdot U^{(\mathrm{S})}$ (21)
generates the LF state from initialized qubits. $\mathrm{QFT}$ on the right-hand side of Eq. (21) can be replaced by $\mathrm{QFT}^{\dagger}.$ As seen in Figs. 3(b) and (c), a discrete SF having a larger decay rate corresponds to a discrete LF having a wider peak. In fact, from Eq. (19), the discrete LF behaves near the origin $(|j|\ll N)$ approximately as $L_{j}(n_{q},a)\propto 1/(j^{2}+\gamma_{L}(a)^{2}),$ where $\gamma_{L}(a)\equiv(N/\pi)\sinh(a/2)$ is a monotonically increasing function of $a.$ In what follows, we denote $a$ for $L(n_{q},a)$ simply as the decay rate of the discrete LF since no confusion would occur. The overlap between two LFs is given as a simple expression, as derived in Appendix G.2. The circuit depth of $U^{(\mathrm{L})}$ implemented as Eq. (21) is, from Eq. (18),

### v3

@@ I Introduction
To this end, we design the main encoding technique by starting from the probabilistic operation [10, 11] for an LC of discrete Lorentzian functions (LFs), to which we apply the quantum amplitude amplification (QAA) technique [12, 13] to render the encoding deterministic. As is demonstrated later, expanding a target function in LFs is favorable for achieving efficient circuit implementation. Our encoding techniques possess applicability to diverse fields of quantum computation. Amongst them, quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] is a promising one, where an initial state can be constructed from one-electron molecular orbitals (MOs) [25, 26], typically expressed as LCs of localized orbitals. We provide resource estimation for such calculations. To demonstrate the efficiency and practical usefulness of our scheme, we perform it on the cloud quantum computing platform commercially provided by IBM [27]. Our encoding scheme is for a dense quantum state; that is, it treats a target function that has basically a nonzero amplitude of each computational basis. For encoding schemes for a sparse quantum state, see, e.g., Refs [28, 29, 30] and references therein.
@@ II.1.1 Probabilistic operation of an LC of unitaries and amplitude reduction
When the original technique of QAA [12, 13] is applied to $\mathcal{C}_{\mathrm{prob}},$ the success probability is made close to 1. The amplified value is, however, not necessarily equal to 1, depending on the original success probability. This fact motivates us to define another circuit $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ for the probabilistic operation of $\mathcal{A}.$ This circuit differs from $\mathcal{C}_{\mathrm{prob}}$ only in that the new one uses an extra ancilla, as depicted in Fig. 1(b). The extra ancilla is accompanied by the $y$ rotation gate $R(2\theta_{\mathrm{AR}})=\exp(-i\theta_{\mathrm{AR}}\sigma_{y})$ with an angle parameter $\theta_{\mathrm{AR}}.$ Although $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ involves $\mathcal{C}_{\mathrm{prob}}$ as its part separated from the extra ancilla, they will be coupled for QAA later. The state $|\Psi_{\mathrm{AR}}\rangle$ of the whole system immediately before the measurement is obviously the tensor product of $|\Psi\rangle$ in Eq. (1) and the extra-ancilla state $\cos\theta_{\mathrm{AR}}|0\rangle+\sin\theta_{\mathrm{AR}}|1\rangle.$ We regard only $|0\rangle_{n_{\mathrm{A}}}\otimes|0\rangle$ to be the success state, even though $|0\rangle_{n_{\mathrm{A}}}\otimes|1\rangle$ also leads to $|\psi_{\mathrm{lc}}\rangle.$ The success probability is thus given by
@@
Figure 1: Each of circuits (a) $\mathcal{C}_{\mathrm{prob}}$ and (b) $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ is for probabilistic operation of an LC $\mathcal{A}$ of unitaries. The explicit construction of the central part $\mathcal{C}$ in $\mathcal{C}_{\mathrm{prob}}$ is described in the original paper [11]. $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ implements the amplitude reduction for the success state by introducing an extra ancilla in addition to those in $\mathcal{C}_{\mathrm{prob}}.$ (c) Each circle represents the weight $W_{\mathrm{f}}$ of failure state in the resultant state of QARA with a QAE error $\Delta w$ as a function of the true weight $w$ of success state in the original state. Each cross represents $W_{\mathrm{f}}$ for QAA-only with a QAE error. Horizontal lines represent the approximate maximum errors in Eq. (6) for the values of $\Delta w/w.$
@@ II.1.2 Determinization via amplitude reduction and amplification
These observations indicate that $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ with $\theta_{\mathrm{AR,opt}}$ and the subsequent QAA render the nonunitary operation of $\mathcal{A}$ deterministic regardless of $w.$ We refer to this treatment as the determinization of the probabilistic operation and denote it by the quantum amplitude reduction and amplification (QARA) technique. While the expected operation number for the determinized circuit does not differ largely than for the probabilistic circuit, the striking feature of the former is that the omission of measurement is justified at the algorithmic level. This helps to develop an efficient algorithm without paying attention to measurement overhead, readout errors, and reset of ancillae. Since $w$ is on the order of $1/n_{\mathrm{loc}},$ we have $m_{\mathrm{opt}}=\mathcal{O}(\sqrt{n_{\mathrm{loc}}})$ from Eq. (4). In particular, for a case where $1/4\leq w<1,$ we have $m_{\mathrm{opt}}=1.$ Nishi et al. [41] have proposed a similar technique for determinizing the QAA for probabilistic imaginary-time evolution (PITE)[18, 20, 42, 22] steps. The quadratic speedup of PITE [43] with determinization will be a good scheme for nonvariational energy minimization on fault-tolerant quantum computers.
@@ II.1.3 Error analysis for inaccurate QAE
Let us consider a practical situation where the observer obtains an erroneous estimation $w^{(\mathrm{est})}=w+\Delta w$ of the weight $w$ of the original success state. We do not introduce any assumption about where the error $\Delta w$ has come from. It is noted here that a statistical error can intrude into the estimation even when the quantum computation is performed fault-tolerantly due to a finite number of measurements. The circuit parameter $\theta_{\mathrm{AR,opt}}$ for amplitude reduction and the repetition number $m_{\mathrm{opt}}$ for amplitude amplification can now involve errors. We demonstrate in Appendix A that QARA allows us to derive a simple relation between the error in the QAE process and that in the determinization. When we approach the limit of $w\to 0$ while keeping $\Delta w/w$ finite, an approximate maximum of the weight of the failure state after the completion of the QARA process is given by
@@
Figure 1(c) shows the weight $W_{\mathrm{f}}$ of the failure state after the QARA process for $\Delta w/w=0.1,0.04,$ and $0.01$ as a function of $w.$ $W_{\mathrm{f}}$ after the QAA-only process (without amplitude reduction) is also shown in the figure. While the values of $W_{\mathrm{f}}$ from QAA-only are seen to scatter for each of $\Delta w/w,$ those from QARA exhibit much smaller variations over the wide range of $w.$ Furthermore, $\epsilon_{\mathrm{QARA}}$ approximates $W_{\mathrm{f}}$ well. These observations suggest that QARA is an error-controllable method for determinizing a probabilistic operation.
@@ II.2.1 Setup
Let us consider a case where an $n_{q}$-qubit data register is available and we are provided with $n_{\mathrm{loc}}$ real functions $\{f_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}$ localized at the origin as an expansion basis set in one-dimensional (1D) space. For the equidistant grid points $x_{j}\equiv j\Delta x\ (j=0,\dots,2^{n_{q}}-1)$ of a spacing $\Delta x$ on a range $[0,N\Delta x],$ we want to encode a normalized LC of the displaced basis functions on the data register as
@@
where $N\equiv 2^{n_{q}}.$ $|j\rangle_{n_{q}}$ is the computational basis for the data register. $k_{\mathrm{c}\ell}$ is the integer coordinate $(0\leq k_{\mathrm{c}\ell}\leq N-1)$ of the center of the displaced $\ell$th basis function, where the subscript c is for ‘center’. $d_{\ell}$ is the real coefficient for the LC. We assume that the basis functions are normalized over the range, that is, $\sum_{j=0}^{N-1}f_{\ell}(x_{j})^{2}=1$ for each $\ell,$ and the displaced ones are not necessarily orthogonal to each other. We further assume that the $n_{q}$-qubit unitary oracle for generating each basis function centered at the origin is known: $U_{\mathrm{orig}}^{(\ell)}|0\rangle_{n_{q}}=\sum_{j=0}^{N-1}f_{\ell}(j\Delta x)|j\rangle_{n_{q}}$ for each $\ell.$ The assumption for the center to be the origin is in fact not necessary since we can cancel the discrepancy by adjusting the parameters in the phase shift unitaries introduced below. We adopt the definition of oracle $U_{\mathrm{orig}}^{(\ell)},$ however, for the discussion to be simple.
@@ II.2.2 Displaced single basis function
for each $\ell.$ To this end, we define the phase shift unitary $U_{\mathrm{shift}}(k)$ for an integer $k$ that acts on a computational basis diagonally as
@@
This operation can be implemented as separate single-qubit unitaries, as explained in Appendix B. With this and the quantum Fourier transform (QFT) operation $\mathcal{F}_{\mathrm{quant}}$ [45], the operator
$T(k)\equiv\mathcal{F}_{\mathrm{quant}}\cdot U_{\mathrm{shift}}(k)\cdot\mathcal{F}_{\mathrm{quant}}^{\dagger}$ (10)
@@ II.2.3 LC of displaced basis functions
$\widetilde{U}^{(\ell)}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathcal{F}_{\mathrm{quant}}^{\dagger}\cdot U_{\mathrm{orig}}^{(\ell)}$ (11)
with $U_{\mathrm{shift}}^{(\ell)}\equiv U_{\mathrm{shift}}(k_{\mathrm{c}\ell})$ for each $\ell.$ With the LC $\mathcal{L}\equiv\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}\widetilde{U}^{(\ell)}$ of them, the desired state in Eq. (7) is written as $|\psi_{\mathrm{lc}}\rangle=\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}|\widetilde{f}_{\ell}\rangle=\mathcal{F}_{\mathrm{quant}}\cdot\mathcal{L}|0\rangle_{n_{q}}.$ $\mathcal{L},$ which is nonunitary in general, admits the probabilistic implementation as a special case of $\mathcal{C}_{\mathrm{prob}}$ in Fig. 1(a). Specifically, the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the data register and
@@
ancillae encodes $|\psi_{\mathrm{lc}}\rangle$ probabilistically. The central part, $\mathcal{C}_{\mathcal{L}},$ implements $\mathcal{L}$ probabilistically by recognizing the $n_{\mathrm{A}}$-bit integer $\ell$ specified by an ancillary state $|\ell\rangle_{n_{\mathrm{A}}}.$ For details, see the original paper [11], which describes the implementation of a generic LC of unitaries. It is noted here that a slight modification to the circuit in the original paper exists in $\mathcal{C}_{\mathcal{L}}.$ That is, since $\mathcal{F}_{\mathrm{quant}}^{\dagger}$ in $\widetilde{U}^{(\ell)}$ does not refer to $\ell$ [see Eq. (11)], all the controlled $\mathcal{F}_{\mathrm{quant}}^{\dagger}$ have been replaced by the uncontrolled $\mathcal{F}_{\mathrm{quant}}^{\dagger},$ as seen in Fig. 2(b). Since the $\mathcal{F}_{\mathrm{quant}}$ operation after $\mathcal{C}_{\mathcal{L}}$ does not act on the ancillae, it is also possible to perform the measurement, observe the success state, and then operate $\mathcal{F}_{\mathrm{quant}}$ on the data register.
Let us estimate the depth of $\mathcal{C}_{\mathrm{lc}}$ by assuming that of $U_{\mathrm{orig}}^{(\ell)}$ is independent of $\ell.$ Since $\mathrm{depth}(\mathcal{F}_{\mathrm{quant}})=\mathcal{O}(n_{q})$ [46, 47] and $\mathrm{depth}(U_{\mathrm{shift}}^{(\ell)})=\mathcal{O}(1)$ for each $\ell,$ we find
@@ II.3.2 Discrete Lorentzian functions
$|L;a,0\rangle_{n_{q}}$ is symmetric around $j=N/2.$ As demonstrated in Appendix G.1, the SF and LF states with a common decay rate at the origin are related with each other via QFT: $|L;a,0\rangle_{n_{q}}=\mathcal{F}_{\mathrm{quant}}|S;a,0\rangle_{n_{q}}.$ This means that the unitary
$U^{(\mathrm{L})}\equiv\mathcal{F}_{\mathrm{quant}}\cdot U^{(\mathrm{S})}$ (21)
generates the LF state from initialized qubits. $\mathcal{F}_{\mathrm{quant}}$ on the right-hand side of Eq. (21) can be replaced by $\mathcal{F}_{\mathrm{quant}}^{\dagger}.$ As seen in Figs. 3(b) and (c), a discrete SF having a larger decay rate corresponds to a discrete LF having a wider peak. In fact, from Eq. (19), the discrete LF behaves near the origin $(|j|\ll N)$ approximately as $L_{j}(n_{q},a)\propto 1/(j^{2}+\gamma_{L}(a)^{2}),$ where $\gamma_{L}(a)\equiv(N/\pi)\sinh(a/2)$ is a monotonically increasing function of $a.$ In what follows, we denote $a$ for $L(n_{q},a)$ simply as the decay rate of the discrete LF since no confusion would occur. The overlap between two LFs is given as a simple expression, as derived in Appendix G.2. The circuit depth of $U^{(\mathrm{L})}$ implemented as Eq. (21) is, from Eq. (18),